Find the margin of error (E) of a confidence estimate interval, when the confidence level is 95%, the sample size is 125, and the standard deviation is 4.
Answer by Boreal(15235)

E=z(0.975)*sd/sqrt(n)
=1.96*4/sqrt(125)=0.70